Advanced Photon Source

An Office of Science National User Facility

APS-EFOG Meeting Minutes - 03/20/2017

Attendees: Bruno Fieramosca, Nena Moonier, Clay White, Shane Flood, Wendy VanWingeren, and Patti Pedergnana

Shane Flood

  • At S24 and S26 the ID setpoint was not returning to its saved value.
  • S19 had the IDD shutter switches reset.
  • S5 had their IDB shutter cylinders rebuilt.
  • 2-ID-D needs an EEI of equipment that has been here in the past.  Bruno is going to perform the inspection.

Clay White

  • The cant angle work at 4-ID has been postponed.  Only prep work for the work will be done in the upcoming shutdown.
  • 8-ID is going to turn the monochromator 180 degrees.
  • The 9-ID mini-hutches may be taken apart this shutdown and temporarily moved to S29 area for storage.  They will be installed at 2-ID permanently in the future.
  • Tunnels will close 5/16-5/17.
  • There will be switchgear and LTHW maintenance this shutdown but there were no areas mentioned at the shutdown planning meeting.
  • Fast valves are going to be removed at 1-BM’s and 6-ID’s Front Ends.
  • Pneumatic shutter solenoids at 3-ID, 4-ID, and 6-ID are going to be replaced this shutdown.

Wendy VanWingeren

  • 32-ID may run radioactive samples on Friday.  The samples are coming from Australia but as of now do not have proper containment.
  • S20 got a NIST standard containing Uranium and Thorium that needs to go to bldg. 200 to be loaded.
  • On Thursday, Martin McBriarty is running radioactive samples at 13-ID-E.